OpenStudy (anonymous):

Find a cubic function with the given zeros: sqrt2, -sqrt2, -2 :)

OpenStudy (anonymous):

general form of cubic eq.- \[x^{3}-(\alpha +\beta+ \gamma)x^{2}+(\alpha \beta+\beta \gamma+\alpha \gamma)x-\alpha \beta \gamma=0\]where alpha beta and gamma are the zeroes. Substitute the values of zeroes.

OpenStudy (anonymous):

or why dont you just multiply out (x-a)(x-b)(x-c) where a,b,c are roots..

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

(x-sqrt2)(x+sqrt2)(x+2)?

OpenStudy (anonymous):

\[x^{3}-(\sqrt{2}-\sqrt{2}-2)x^{2}+(-\sqrt{2}\sqrt{2}+2\sqrt{2}-2\sqrt{2})x-2\sqrt{2}\sqrt{2}=0\] \[x^{3}-(-2)x^{2}+(-2)x-2*2=0\] \[x^{3}+2x^{2}-2x-4=0\]

OpenStudy (anonymous):

So, x^2-2 (x+2) x^3 +2x^2 -2x -4?

OpenStudy (anonymous):

Oh. Yeah!